OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 14 inches and a height of 6 inches. Pyramid B has a volume of 3136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (anonymous):

Which of these countries has the greatest population density? Which has the least? How can you tell?

OpenStudy (anonymous):

@RaphaelFilgueiras

OpenStudy (anonymous):

@lolly

OpenStudy (anonymous):

For the first question that you posted use the formula for volume of square pyramid \(v=b^2 \frac{h}{3}\) to calculate the volume of pyramid A. then compare it to the volume of pyramid B. For the next one: \(population\ density= \frac{population}{area}\) then compare the population density of the three countries

OpenStudy (anonymous):

for the volume v=volume b=base h=height

OpenStudy (anonymous):

b-14 h-6 Pyramid B has a volume of 3136 cubic inches.

OpenStudy (anonymous):

using the formula, pyramid A will have a volume of:?

OpenStudy (anonymous):

could you help me i really don't understand

OpenStudy (anonymous):

@hlee13

OpenStudy (anonymous):

solving for the volume of pyramid A: b=14 h=6 \(v=b^2 \frac{h}{3}\\v=(14)^2 (\frac{6}{3})\\v=?\)

OpenStudy (anonymous):

392

OpenStudy (anonymous):

@Chibi_Robo3

OpenStudy (anonymous):

the question asks "how many time b is bigger than a" so get the ratio of both volumes b:a 3136 / 392= ?

OpenStudy (anonymous):

so the volume of pyramid B will be 8 x bigger than pyramid A

OpenStudy (anonymous):

Ok can ya help me with 2 @Chibi_Robo3

OpenStudy (anonymous):

use the formula that i gave you above for the second one do it for each countries and tell me what you got this one is pretty straightforward, so i believe you can do it :)

OpenStudy (anonymous):

U.S--29.11500002135721

OpenStudy (anonymous):

Denmark--0.1247237207975854 Ghana--0.0772061145188248

OpenStudy (anonymous):

@Chibi_Robo3

OpenStudy (anonymous):

US is right, but for the two countries, i think you solved it in a wrong way

OpenStudy (anonymous):

try it again

OpenStudy (anonymous):

could you help me i did them the same way i did US

OpenStudy (anonymous):

populatin/area?

OpenStudy (anonymous):

i think you did the other way, instead of "population/area", you used "area/population"

OpenStudy (anonymous):

oh ok let me redo them

OpenStudy (anonymous):

Denmark--8.017721036585366 Ghana--12.95234200338077

OpenStudy (anonymous):

@Chibi_Robo3

OpenStudy (anonymous):

us: 29.11 denmark : 124.72 ghana: 77.21 from these values, can you answer the questions now?

OpenStudy (anonymous):

Denmark is the density then Ghana the us is the least density

OpenStudy (anonymous):

right ;p